OpenStudy (anonymous):

What is the area of a sector with radius equal to 8 and measure of arc equal to 300°?

hartnn (hartnn):

Area of sector is \(\Large A = (1/2) r^2\theta \) r =8 but theta there should be in radians. so, can you first convert 300 degrees to radians ?? (hint : \(\Large 180^o= \pi ^r\))